OpenStudy (anonymous):

how do you solve \sqrt{56-r}=r

OpenStudy (anonymous):

56-r=r² => r²+r-56 =0 => delta = 1-4(-56) = 1+ 224=225>0 r1=(-1-sqrt(225))/2=(-1-15)/2=-4 or r2=(-1+sqrt(225))/2=(-1+15)/2=7 r=

OpenStudy (anonymous):

ummm i think its the 1st one

OpenStudy (anonymous):

No, sorry another: \[\sqrt{56-r} =r \implies r^2 = 56 - r \implies r^2 + r - 56 = 0\] Therefore (r+8)(r-7) = 0 => r = -8, 7 REMEMBER Squaring can produce spurious roots, but in this case, we check both and they work, so we are fine!

OpenStudy (anonymous):

oh ok thank u